# Fine-Tuning Large Audio-Language Models with LoRA for Precise Temporal Localization of Prolonged Exposure Therapy Elements thanks: We thank National Science Foundation under Grant # 2326144

### v1

@@ Abstract
Prolonged Exposure (PE) therapy is an effective treatment for post-traumatic stress disorder (PTSD), but evaluating therapist fidelity remains labor-intensive due to the need for manual review of session recordings. We present a method for the automatic temporal localization of key PE fidelity elements—identifying their start and stop times—directly from session audio and transcripts. Our approach fine-tunes a large pre-trained audio-language model, Qwen2-Audio, using Low-Rank Adaptation (LoRA) to process focused 30-second windows of audio-transcript input. Fidelity labels for three core protocol phases—therapist orientation (P1), imaginal exposure (P2), and post-imaginal processing (P3)—are generated via LLM-based prompting and verified by trained raters. The model is trained to predict normalized boundary offsets using soft supervision guided by task-specific prompts. On a dataset of 313 real PE sessions, our best configuration (LoRA rank 8, 30s windows) achieves a mean absolute error (MAE) of 5.3 seconds across tasks. We further analyze the effects of window size and LoRA rank, highlighting the importance of context granularity and model adaptation. This work introduces a scalable framework for fidelity tracking in PE therapy, with potential to support clinician training, supervision, and quality assurance.
@@ II RELATED WORK
Automated analysis of psychotherapy sessions aims to provide scalable feedback, support training, and enhance therapeutic outcomes. Our work leverages pre-trained audio-language models to temporally localize psychotherapy fidelity metrics in session audio using multimodal inputs, using soft-supervision, where LLM-proposed timestamps are clinically verified.
@@ II-F Research Contribution
While existing research has advanced psychotherapy analysis, multimodal learning, and temporal event detection, our work addresses temporal localization of psychotherapy fidelity metrics using pre-trained models, researcher-driven prompts, and soft-supervision—laying the foundation for fine-grained, scalable fidelity assessment in clinical contexts.
@@ III DATASET
We begin with a dataset comprises 318 PE therapy sessions (see Table II) recorded at Emory University. Each 60-90 minute session contains a therapist-client interaction. All data is handled within IRB-approved environments with no PII leaving secure storage. We annotate PE therapy protocol elements: orientation (P1), imaginal exposure (P2), and processing (P3). As shown in Figure 1, P1, P2, and P3 typically occur sequentially across a 60–90 minute PE session, motivating our timestamp prediction task. To minimize manual effort, we implemented a zero-shot LLM pipeline that processes transcripts to produce JSON-formatted protocol presence labels with timestamp boundaries. To validate these labels while ensuring scalability, 10% (30) of the sessions were reviewed by trained raters. These raters were specifically trained to identify the presence and temporal boundaries of the key therapeutic elements, a process which yielded 94.4% timestamp accuracy (within 5-10 seconds) and 97.7% label prediction accuracy. This approach demonstrates the feasibility of our framework without requiring costly expert clinician time for the verification step. Audio files were downsampled from 44.1-48 kHz to 16 kHz, normalized, and converted to .wav format. Transcripts were generated using Amazon HealthScribe with sentence-level timestamps and speaker labels. After filtering sessions exceeding 1.5 hours (limitation of transcription service) or having alignment issues, we retained 313 recordings split into 219 training, 46 validation, and 48 testing sessions, with balanced distribution of therapy phases across splits.
@@
Number of files 318
Total duration 342.11 hours
Min duration 00:19:30 (1170.75 s)
Max duration 01:31:12 (5472.38 s)
Average duration 01:04:32.9 (3872.89 s)
Std. deviation 00:15:38.5 (938.55 s)
@@ IV-B Dataset and Soft-Supervision Labeling
Our processed dataset consists of 313 PE therapy sessions, preprocessed as described in Section III, and transcribed using Amazon HealthScribe, which provides sentence-level timestamps and speaker attribution. Ground-truth start and stop times for fidelity metrics P1, P2, and P3 were initially generated via soft supervision using Claude Sonnet 3.5, prompted to return timestamps in seconds as illustrated in Table I. These LLM-generated annotations were subsequently reviewed and corrected by trained trained raters to create verified ground-truth labels. All audio recordings were downsampled to 16kHz and split at the session level into train, validation, and test sets.
@@ IV-C Model Architecture and Fine-tuning
We fine-tune Qwen2-Audio-7B-Instruct [9], a large audio-language model that processes interleaved audio and text inputs through an audio encoder and a language model backbone. For each trained rater-verified boundary time, we construct focused input windows of fixed durations (30, 60, or 120 sec) containing both audio and aligned transcript segments. To ensure model robustness and enable data augmentation, the true boundary is randomly positioned within each window by normalizing its location to the $[0,1]$ interval and adjusting the window center accordingly. The training target is the normalized offset of the boundary relative to the window start. Each input is prepended with a task-specific prompt, such as: “The following audio and transcript segment is focused around the START of ’Was Prolonged Exposure done in the session?’. Identify the normalized offset (between 0.0 and 1.0) of this precise START within the given segment.” We use QLoRA [11] for memory-efficient fine-tuning, applying 4-bit NormalFloat (NF4) quantization with bfloat16 computation, and LoRA [10] to the LLM component with ranks $r\in{2,4,8}$, scaling factor $\alpha=2r$, and dropout of 0.1. A regression head receives the final hidden state of the last non-padding token and predicts the normalized offset via a LayerNorm, two linear layers with ReLU activation, and a final Sigmoid function. All experiments were conducted using three random seeds (42, 78, and 123), and results are reported as mean $\pm$ standard deviation across these runs.
@@ V EXPERIMENTAL SETUP
We implemented the temporal localization framework described in Section IV using Qwen2-Audio-7B-Instruct [9]. Experiments were conducted on a dataset of 313 PE therapy sessions, split into 219 train, 46 validation, and 48 test files as detailed in Section III. To assess the influence of context length and fine-tuning strategies, we varied (1) input audio-transcript window durations (30s, 60s, and 120s), and (2) LoRA ranks $r\in\{2,4,8\}$. These configurations were compared against a “Head Only” baseline where only the final regression layer was trained. All other QLoRA and training parameters (optimizer, learning rate, epochs, etc.) followed the specifications described in Section IV. Performance was evaluated using MAE in seconds for PE elements P1, P2, and P3.

### v2

@@ Abstract
Prolonged Exposure (PE) therapy is an effective treatment for post-traumatic stress disorder (PTSD), but evaluating therapist fidelity remains labor-intensive due to the need for manual review of session recordings. We present a method for the automatic temporal localization of key PE fidelity elements—identifying their start and stop times—directly from session audio and transcripts. Our approach fine-tunes a large pre-trained audio-language model, Qwen2-Audio, using Low-Rank Adaptation (LoRA) to process focused 30-second windows of audio-transcript input. Fidelity labels for three core protocol phases—therapist orientation (P1), imaginal exposure (P2), and post-imaginal processing (P3)—are generated via LLM-based prompting and verified by trained raters. The model is trained to predict normalized boundary offsets using soft supervision guided by task-specific prompts. On a dataset of 308 real PE sessions, our best configuration (LoRA rank 8, 30s windows) achieves a mean absolute error (MAE) of 5.3 seconds across tasks. We further analyze the effects of window size and LoRA rank, highlighting the importance of context granularity and model adaptation. This work introduces a scalable framework for fidelity tracking in PE therapy, with potential to support clinician training, supervision, and quality assurance.
@@ II RELATED WORK
Automated psychotherapy analysis enables scalable feedback, clinician training, and improved outcomes. We use pre-trained audio-language models with multimodal inputs to temporally localize fidelity metrics, using soft supervision with LLM-generated timestamps clinically verified for accuracy.
@@ II-F Research Contribution
While existing research has advanced psychotherapy analysis, multimodal learning, and temporal event detection, our work addresses temporal localization of psychotherapy fidelity metrics using pre-trained models, researcher-driven prompts, and soft-supervision—laying the foundation for fine-grained, scalable fidelity assessment in clinical contexts. Our work builds on recent advances in synthetic PE therapy datasets [32], fidelity evaluation frameworks [33], and empathy modeling in PTSD dialogue systems [34].
@@ III DATASET
We use a dataset of 318 PE therapy sessions (see Table II) recorded at Emory University, each lasting 60–90 minutes and capturing therapist-client interactions. All data were handled within IRB-approved environments, with no PII leaving secure storage. We annotated key protocol elements—orientation (P1), imaginal exposure (P2), and processing (P3)—which typically occur sequentially during a session (Figure 1), motivating our timestamp prediction task. To minimize manual labeling, we developed a zero-shot LLM pipeline that extracts protocol presence and timestamp boundaries from transcripts in JSON format. To ensure scalable validation, 10% (30 sessions) were reviewed by trained raters, yielding 94.4% timestamp accuracy (within 5–10 seconds) and 97.7% label accuracy. This validated our approach without requiring extensive expert clinician time. Audio was downsampled from 44.1–48 kHz to 16 kHz, normalized, and converted to .wav. Transcripts with sentence-level timestamps and speaker labels were generated using Amazon HealthScribe. After excluding sessions over 1.5 hours (due to transcription limits), those with alignment issues, or lacking valid labels for all protocol elements, we retained 308 recordings. The dataset was split into 216 training, 45 validation, and 47 test sessions, with balanced representation of therapy phases.
@@
Number of sessions 308
Total duration 337.76 hours
Min duration 00:19:31 (1170.75 s)
Max duration 01:29:54 (5393.86 s)
Average duration 01:05:48 (3947.84 s)
Std. deviation 00:15:36 (935.53 s)
@@ IV-B Dataset and Soft-Supervision Labeling
Our processed dataset consists of 308 PE therapy sessions, preprocessed as described in Section III, and transcribed using Amazon HealthScribe, which provides sentence-level timestamps and speaker attribution. Ground-truth start and stop times for fidelity metrics P1, P2, and P3 were initially generated via soft supervision using Claude Sonnet 3.5, prompted to return timestamps in seconds as illustrated in Table I. These LLM-generated annotations were subsequently reviewed and corrected by trained trained raters to create verified ground-truth labels. All audio recordings were downsampled to 16kHz and split at the session level into train, validation, and test sets.
@@ IV-C Model Architecture and Fine-tuning
We fine-tune Qwen2-Audio-7B-Instruct [9], a large audio-language model that processes interleaved audio and text inputs through an audio encoder and a language model backbone. For each trained rater-verified boundary time, we construct focused input windows of fixed durations (30, 60, or 120 sec) containing both audio and aligned transcript segments. To ensure model robustness and enable data augmentation, the true boundary is randomly positioned within each window by normalizing its location to the $[0,1]$ interval and adjusting the window center accordingly. The training target is the normalized offset of the boundary relative to the window start. Each input is prepended with a task-specific prompt, such as: “The following audio and transcript segment is focused around the START of ’Was Prolonged Exposure done in the session?’. Identify the normalized offset (between 0.0 and 1.0) of this precise START within the given segment.” We use QLoRA [11] for memory-efficient fine-tuning, applying 4-bit NormalFloat (NF4) quantization with bfloat16 computation, and LoRA [10] to the LLM component with ranks $r\in{2,4,8}$, scaling factor $\alpha=2r$, and dropout of 0.1. A regression head uses the final hidden state of the last non-padding token to predict the normalized offset via LayerNorm, two linear layers with ReLU, and a final Sigmoid. Experiments used three random seeds (42, 78, 123), with results reported as mean $\pm$ standard deviation across these runs.
@@ V EXPERIMENTAL SETUP
We implemented the temporal localization framework described in Section IV using Qwen2-Audio-7B-Instruct [9]. Experiments were conducted on a dataset of 308 PE therapy sessions, split into 216 train, 45 validation, and 47 test sessions as detailed in Section III. To assess the influence of context length and fine-tuning strategies, we varied (1) input audio-transcript window durations (30s, 60s, and 120s), and (2) LoRA ranks $r\in\{2,4,8\}$. These configurations were compared against a “Head Only” baseline where only the final regression layer was trained. All other QLoRA and training parameters (optimizer, learning rate, epochs, etc.) followed the specifications described in Section IV. Performance was evaluated using MAE (sec) for PE elements P1, P2, and P3.
@@ VII CONCLUSION
VIII ACKNOWLEDGMENT
This work is supported by the National Science Foundation (NSF) under Grant No. 2326144. Any opinions, findings, & conclusions or recommendations expressed in this work are those of the author(s) and do not necessarily reflect the views of the NSF.